OpenStudy (anonymous):

A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air? (2 points)

OpenStudy (radar):

Let x = aircraft speed in still air. x+40 = aircraft speed with 40 MPH tail wind. x-40 = aircraft speed with 40 MPH headwind. D= distance Denver to Orlando D = 5(x+40) D = 6(x-40) D=D 5x + 200 = 6x-240 x=200+240= 440 mph in still wind. (2 points)